OpenStudy (marysammonss):

Solve the equation for x. -6x + 8x = -46 A) -23 B) 3.2 C) 12 D) 16

OpenStudy (sprattfield):

wait

OpenStudy (alivejeremy):

-23

KjSaif (kjsaif):

No direct answers bro.........@alivejeremy xD

OpenStudy (alivejeremy):

I'm bout to explain give me a sec gg lol

OpenStudy (sprattfield):

Yeah no direct answers please

OpenStudy (alivejeremy):

First, simplify both sides of the equation.

OpenStudy (alivejeremy):

Then divide both sides by 2.

OpenStudy (sprattfield):

Move all terms that don't contin x to the right side and solve

KjSaif (kjsaif):

−6x+8x=−46 Step 1: Simplify both sides of the equation. −6x+8x=−46 Simplify: (Show steps) 2x=−46 Step 2: Divide both sides by 2. 2x\2=−46/2 x=−23

OpenStudy (alivejeremy):

and your answer

OpenStudy (alivejeremy):

x=−23

OpenStudy (marysammonss):

lol thankssss i get it now :)

OpenStudy (sprattfield):

we are here for you

OpenStudy (alivejeremy):

np.

KjSaif (kjsaif):

:)

OpenStudy (alivejeremy):

Ayeee everyone got a medal lol.